OpenStudy (anonymous):

Solve for b2 in A = 1/2h(b1+b2), if A = 16, h = 4, and b1 = 3.

OpenStudy (anonymous):

16=½*4(3+b2) 16=2(3+b2) 16=6+2b2 16-6=2b2 10=2b2 b2=5

OpenStudy (anonymous):

thanx